OpenStudy (anonymous):

Verify the following: Solution: y=x^2 differential eq: xy'-2y=x^3•e^x

OpenStudy (anonymous):

x 2x -2 x^2= x^2* e^x 2x^2-2x^2=x^2 e^x 0=x^2 e^x no! I am not sure

myininaya (myininaya):

y'=2x plug in: \[x(2x)-2(x^2)=x^3e^x\] \[2x^2-2x^2=x^3e^x\] yep it is not

myininaya (myininaya):

unless x=0

OpenStudy (anonymous):

y=x^2 y'=2x now, xy'-2y = x (2x)-2(x^2) = 2x^2-2x^2=0

OpenStudy (anonymous):

Yeah, so it's not verifiable.

OpenStudy (anonymous):

yes , it is not verified

OpenStudy (across):

It's a linear differential equation: \[xy'-2y=x^3e^x\]Arrange it into a solvable form: \[y'-(2/x)y=x^2e^x\]Find the integrating factor: \[\mu(x)=e^{\int\limits-(2/x)dx}=e^{-2\ln|x|}=e^{\ln|x|^{-2}}=x^{-2}\]Multiply everything by this factor: \[(x^{-2}y)'=e^x\]Integrate both sides and solve for y: \[y=x^{2}e^{x}+c\]You can see that the solutions don't match.